5) A brand name as a 20% recognition rate. Assume the owner of the brand wants to verify that rate by beginning with a smallA brand name has 20% recognition rate.
Answer #1

X ~ B(4, 0.20)

Hence,

a) P(X = 3)

= (4) (0.2013(0.80)**

= 0.0256

b) P(All 4) = (0.20)4 = 0.0016

c) P(Atleast 3) = P(3) + P(4) = 0.0256 + 0.0016 = 0.0272

d) Yes, atleast 3 is unusually high number because its probability is less than 0.05.